OpenStudy (krystal):

50000-x x is the amount of in $ ann invested in the account that pays 2.5% annual interset

OpenStudy (anonymous):

ok what are you trying to solve for?

OpenStudy (krystal):

she invests 50,000 in two interst bearing accounts one pays 2.5 % interst and the second 3.75% at the end of one year the interest ann has earned on these is 1656.25 $ how much she invest in each of the accounts

OpenStudy (anonymous):

ok got it!

OpenStudy (anonymous):

so what in an interest account, each year you earn the % of interest according to your original investment

OpenStudy (anonymous):

so from the first account, she will earn 2.5% of 50,000

OpenStudy (anonymous):

from the second account, she will earn 3.75% of 50,000

OpenStudy (anonymous):

or sorry, bad first assumption

OpenStudy (anonymous):

so we have two equations based on this information:

OpenStudy (anonymous):

both involve amounts x and y

OpenStudy (anonymous):

we know that .025x + .0375y = 1656.25

OpenStudy (anonymous):

that is, the interest earned between the two accounts adds up to 1656.25, and by definition .025 of X is earned, and .0375 of y is earned in that year

OpenStudy (anonymous):

the other thing we know is that the total amount she invests is 50,000, so x + y = 50,000

OpenStudy (anonymous):

now since you have two variables, and two equations, you can use substitution to solve this

OpenStudy (anonymous):

so using the second equation, we know that y = 50,000 - x

OpenStudy (krystal):

i have to give a vebal descritpion of 50000-x

OpenStudy (anonymous):

I'm not sure what you mean by verbal description

OpenStudy (anonymous):

but we know that 50000 - x is the amount invested in one of the banks

OpenStudy (krystal):

it says give a verbal description of what each of the follwoing expression represetn in the context of the problem letting x=2.5 %. 50,000 is the interest in the two accounts together

OpenStudy (krystal):

she earned a total on the investements $1656.25

OpenStudy (anonymous):

right. so what is the expression? 50,000 - x is the amount of money invested in the second bank, if we assume x is the amount invested in the 2.5% account

OpenStudy (anonymous):

we know that because the total amounts are equal to 50000

OpenStudy (anonymous):

so x+y = 50000

OpenStudy (anonymous):

y = 50000 - x

OpenStudy (anonymous):

and if x is the amount in the first account, 50000 - y is the amount in the 3.75% account

OpenStudy (anonymous):

is that helping?

OpenStudy (krystal):

i still dont get it but i think i understand ur saying if both x and y equally to 50000 than if you take away .025 away from that than it will give you the answer correct

OpenStudy (krystal):

i need to know how to write it out

OpenStudy (krystal):

if she has a total at the end of the year of 1656.25 it is asking how much did she invest in both accounts with one being 0.025 and one being 0.0375 annual interest

OpenStudy (anonymous):

Since: 1. x + y = 50000 2. .025x + .0375y = 1656.25 We substitute for y in the second equation, using the first. i.e. y = 50000 -x Now substitute for y .025x + .0375 (50000 - x) = 1656.25. .025x + 1875 - .0375x = 1656.25 -.0125x = 1656.25 - 1875 -.0125x = -218.75 x = -218.75/-.0125 x = 17500

OpenStudy (anonymous):

so looking above, she invested 17,500 in the bank with 2.5% interest rate (also known as .025 interest)

OpenStudy (anonymous):

now that you've solved for one of the values, you can solve for the other (y) by sticking this result back into either equation

OpenStudy (anonymous):

so x+5 = 50000

OpenStudy (anonymous):

*x + y

OpenStudy (anonymous):

since x = 17500, 17500 + y = 50000

OpenStudy (anonymous):

y = 32500

OpenStudy (anonymous):

so she invested 32500 in the second bank

OpenStudy (anonymous):

(the one with .0375 interest)

OpenStudy (anonymous):

now the final part is - check the answer. that is, at the end of the year, do those two values give you the correct total earned interest

OpenStudy (anonymous):

so .025*17500 + 32500*.0375 = 437.50 + 1218.75 = 1656.25

OpenStudy (anonymous):

so we know it's right

OpenStudy (anonymous):

make sense?

OpenStudy (krystal):

kinda